How much Ca OH2 (molecular weight 74) should be added to 200ml of H2O to make an 18%(w/v) solution


1
Expert's answer
2021-02-04T07:04:44-0500

%w/v = 18%

Mass of solute= ?

Volume of solution= 200ml= 0.20L


%w/v = mass of solute/litres of solution x 100%

Mass= 18 x 0.20/100

Mass= 0.036g of Ca(OH)2


Molar mass of Ca(OH)2 is 74g/mol

Mole = mass/molar mass

Mole= 0.036/74 = 4.86 x 10-4 mol